# Finding Dori [Uncaptioned image]: Memorization in Text-to-Image Diffusion Models Is Less Local Than Assumed

### v1

@@ E.6 Iterative Application of NeMo
Table 7: We apply NeMo [14] iteratively such that after each round of pruning, we search for new adversarial embeddings that can still trigger memorization, and then apply NeMo again to prune the newly identified weights. Despite multiple iterations, this process does not completely eliminate memorization, as adversarial embeddings can still uncover residual memorized content. Due to the high computational cost of repeated NeMo applications and searching for adversarial embeddings, we focus our analysis on prompts known to be verbatim memorized. In some cases, after several iterations, NeMo no longer detects any memorization. When this happens, we analyze the outputs generated from the adversarial embeddings that NeMo failed to flag.
Method Iterations $\boldsymbol{\downarrow}\bm{\text{SSCD}_{\text{Orig}}}$ $\boldsymbol{\downarrow}\bm{\text{SSCD}_{\text{Gen}}}$ $\boldsymbol{\downarrow}\bm{D_{\text{SSCD}}}$ $\boldsymbol{\uparrow}\bm{A_{\text{CLIP}}}$
NeMo [14] Adv. Images 1 $0.92\pm 0.03$ $0.94\pm 0.03$ $1.00\pm 0.00$ $0.33±0.02$
2 $0.92\pm 0.03$ $0.94\pm 0.03$ $1.00\pm 0.00$ $0.32±0.02$
3 $0.92\pm 0.03$ $0.95\pm 0.03$ $1.00\pm 0.00$ $0.33±0.01$
4 $0.92\pm 0.03$ $0.95\pm 0.03$ $1.00\pm 0.00$ $0.33±0.02$
5 $0.92\pm 0.03$ $0.95\pm 0.03$ $1.00\pm 0.00$ $0.32±0.02$
NeMo [14] Mitigated Images 1 $0.35\pm 0.19$ $0.34\pm 0.22$ $0.48±0.14$ $0.34±0.02$
2 $0.23\pm 0.09$ $0.23\pm 0.11$ $0.41±0.09$ $0.35±0.02$
3 $0.21\pm 0.09$ $0.20\pm 0.09$ $0.39±0.08$ $0.35±0.02$
4 $0.20\pm 0.07$ $0.19\pm 0.08$ $0.39±0.08$ $0.34±0.02$
5 $0.20\pm 0.07$ $0.19\pm 0.08$ $0.37±0.08$ $0.34±0.02$
E.7 Cost of Successful Memorization Removal with Wanda
@@
The results in Tab. 6 indicate that Wanda might be successful in removing memorized content from the model (low SSCDOrig at sparsity 10%) with limited harm to the alignment between the prompt and the generated images for benign input (high $\bm{A}_{\text{CLIP}}$). However, FID scores appear to increase significantly with pruning (increase from 16.68 to 37.34 for VM samples). We investigate the harm that Wanda with 10% weights pruned causes to the model. The weights pruned by Wanda not only correspond to the memorized image, but also partially encode concepts present in the memorized content. For example, the memorized image in Fig. 1 ( 1) consists also of a concept of woman. We show that in order to fully remove the image from the model, weights responsible for benign concepts, present in memorized data, are negatively affected.
@@
We verify that idea by generating 100 images from a set of 10 prompts, which are paraphrases of the memorized prompts. Then, we compute CLIP similarity between the paraphrases and generated images, $\bm{A}_{\text{Concepts}})$, to capture how the alignment changes with high pruning. The paraphrases are obtained by prompting LLama-3.1-8B-Instruct [11], with the system prompt "You are a paraphrasing engine. Preserve every tag and keyword.", and the user prompt "Write 10 alternative phrasings of:’’’CAPTION’’’. Return only the paraphrasings, no other text. The format should be [’prompt1’, ’prompt2’, ...]". For example, for the prompt "Living in the Light with Ann Graham Lotz" we obtain "Embracing Life in the Radiance of God with Ann Graham Lotz", "A Life of Radiant Faith with Ann Graham Lotz", "Living Life in the Illumination of God with Ann Graham Lotz", "In the Presence of God’s Radiant Light with Ann Graham Lotz", "Faith in the Light of God with Ann Graham Lotz", "Radiant Living with Ann Graham Lotz", "In God’s Illuminating Light with Ann Graham Lotz", "Ann Graham Lotz on Living in God’s Radiant Presence", "Radiant Faith Living with Ann Graham Lotz", "Living Life in God’s Illuminating Light with Ann Graham Lotz". Additionally, we quantify image quality of the concepts by computing FID score (denoted FIDConcept) between 10,000 images generated from the prompts before and after pruning for VM and TM samples.
@@
The results in Tab. 8 show that the concepts associated with the memorized images suffer after mitigation with Wanda. We observe a significant drop from $\bm{A}_{\text{CLIP}}$ of around $0.37$ to $0.33$ for VM, which suggests that the generated images no longer follow the prompt. Moreover, the quality of the generated images degrades. FIDConcepts above 80 for VM and above 90 for TM samples indicates significant harm to the model, corroborated by the last row of Tab. 6. Additionally, we provide qualitative results of damage to concepts in Sec. H.4.
@@
Setting Memorization SSCDOrig $\bm{A}_{\text{Concepts}}$ FIDConcepts
No Mitigation Verbatim $0.90\pm 0.04$ $0.37\pm 0.02$ N/A
Template $0.17\pm 0.09$ $0.36\pm 0.02$ N/A
Wanda + 10% pruned + [Uncaptioned image] Verbatim $0.40\pm 0.13$ $0.33\pm 0.02$ 80.70
Template $0.10\pm 0.04$ $0.33\pm 0.02$ 92.46
Appendix F Additional Details and Experiments on Adversarial Fine-Tuning
F.1 Algorithmic Description
Alg. 2 provides an algorithmic overview of our adversarial fine-tuning mitigation method. As described in the main paper, we begin by generating images from memorized prompts using a mitigation technique that preserves alignment with the prompt while avoiding replication of training data. Alternatively, these images can be generated using a separate diffusion model that has not been trained on the corresponding samples. To preserve the model’s general utility, a second set of non-memorized samples is incorporated during fine-tuning.
@@ F.1 Algorithmic Description
Algorithm 2 Fine-Tuning Diffusion Model to Mitigate Memorization
      Diffusion model $\bm{\epsilon}_{\bm{\theta}}$
@@
      Surrogate images $\mathcal{D}_{\text{surrogate}}$triangleright Images generated with active mitigation
@@
              ${\bm{y}}_{\text{adv}}\leftarrow\text{find\_adv\_embedding}({\bm{x}}_{\text{mem}},{\bm{p}}_{\text{mem}})$triangleright Start from text embedding
@@
              ${\bm{y}}_{\text{adv}}\leftarrow\text{find\_adv\_embedding}({\bm{x}}_{\text{mem}},\text{random})$triangleright Start from random embedding
@@
              $\bm{\epsilon}_{\text{surr}}\sim\mathcal{N}(0,\bm{I})$triangleright Update with surrogate/adversarial sample
@@
              $\bm{\epsilon}_{\text{non-mem}}\sim\mathcal{N}(0,\bm{I})$triangleright Update with non-memorized sample
@@
              $\mathcal{L}_{\text{total}}\leftarrow\mathcal{L}_{\text{adv}}+\mathcal{L}_{\text{non-mem}}$triangleright Aggregate both losses
@@
F.2 Sensitivity Analysis
@@
Tab. 10 repeats the previous analysis, but explores the impact of starting the adversarial optimization from random embeddings instead of training prompt embeddings. We additionally compare the mitigation effect of adversarial embeddings crafted with 50 and 100 optimization steps, respectively.
Setting Adv. Steps Epochs $\boldsymbol{\downarrow}\bm{\text{SSCD}_{\text{Orig}}}$ $\boldsymbol{\downarrow}\bm{\text{SSCD}_{\text{Gen}}}$ $\boldsymbol{\downarrow}\bm{D_{\text{SSCD}}}$ $\boldsymbol{\uparrow}\bm{A_{\text{CLIP}}}$ $\boldsymbol{\downarrow}\bm{\text{FID}}$ $\boldsymbol{\downarrow}\bm{\text{KID}}$

### v2

@@ H.6 Iterative Application of NeMo
Table 9: We apply NeMo Hintersdorf et al. (2024) iteratively such that after each round of pruning, we search for new adversarial embeddings that can still trigger memorization, and then apply NeMo again to prune the newly identified weights. Despite multiple iterations, this process does not completely eliminate memorization, as adversarial embeddings can still uncover residual memorized content. Due to the high computational cost of repeated NeMo applications and searching for adversarial embeddings, we focus our analysis on prompts known to be verbatim memorized. In some cases, after several iterations, NeMo no longer detects any memorization. When this happens, we analyze the outputs generated from the adversarial embeddings that NeMo failed to flag.
[…]
NeMo Hintersdorf et al. (2024) Adv. Images 1 $0.92\pm 0.03$ $0.94\pm 0.03$ $1.00\pm 0.00$ $0.33\pm 0.02$ $0.99$
2 $0.92\pm 0.03$ $0.94\pm 0.03$ $1.00\pm 0.00$ $0.32\pm 0.02$ $1.00$
3 $0.92\pm 0.03$ $0.95\pm 0.03$ $1.00\pm 0.00$ $0.33\pm 0.01$ $0.99$
4 $0.92\pm 0.03$ $0.95\pm 0.03$ $1.00\pm 0.00$ $0.33\pm 0.02$ $0.99$
5 $0.92\pm 0.03$ $0.95\pm 0.03$ $1.00\pm 0.00$ $0.32\pm 0.02$ $0.99$
NeMo Hintersdorf et al. (2024) Mitigated Images 1 $0.35\pm 0.19$ $0.34\pm 0.22$ $0.48\pm 0.14$ $0.34\pm 0.02$ $0.18$
2 $0.23\pm 0.09$ $0.23\pm 0.11$ $0.41\pm 0.09$ $35\pm 0.02 0.0$40.04
3 $0.21\pm 0.09$ $0.20\pm 0.09$ $0.39\pm 0.08$ $35\pm 0.02 0.0$10.01
4 $0.20\pm 0.07$ $0.19\pm 0.08$ $0.39\pm 0.08$ $34\pm 0.02 0.0$00.00
5 $0.20\pm 0.07$ $0.19\pm 0.08$ $0.37\pm 0.08$ $34\pm 0.02 0.0$00.00
H.7 Cost of Successful Memorization Removal with Wanda
@@
The results in Table˜8 indicate that Wanda might be successful in removing memorized content from the model (low SSCDOrig at sparsity 10%) with limited harm to the alignment between the prompt and the generated images for benign input (high $\bm{A}_{\text{CLIP}}$). However, FID scores appear to increase significantly with pruning (increase from 16.68 to 37.34 for VM samples). We investigate the harm that Wanda with 10% weights pruned causes to the model. The weights pruned by Wanda not only correspond to the memorized image, but also partially encode concepts present in the memorized content. For example, the memorized image in Fig.˜1 ( 1) consists also of a concept of woman. We show that in order to fully remove the image from the model, weights responsible for benign concepts, present in memorized data, are negatively affected.
@@
We verify that idea by generating 100 images from a set of 10 prompts, which are paraphrases of the memorized prompts. Then, we compute CLIP similarity between the paraphrases and generated images, $\bm{A}_{\text{Concepts}})$, to capture how the alignment changes with high pruning. The paraphrases are obtained by prompting LLama-3.1-8B-Instruct (Grattafiori et al., 2024), with the system prompt "You are a paraphrasing engine. Preserve every tag and keyword.", and the user prompt "Write 10 alternative phrasings of:’’’CAPTION’’’. Return only the paraphrasings, no other text. The format should be [’prompt1’, ’prompt2’, ...]". For example, for the prompt "Living in the Light with Ann Graham Lotz" we obtain "Embracing Life in the Radiance of God with Ann Graham Lotz", "A Life of Radiant Faith with Ann Graham Lotz", "Living Life in the Illumination of God with Ann Graham Lotz", "In the Presence of God’s Radiant Light with Ann Graham Lotz", "Faith in the Light of God with Ann Graham Lotz", "Radiant Living with Ann Graham Lotz", "In God’s Illuminating Light with Ann Graham Lotz", "Ann Graham Lotz on Living in God’s Radiant Presence", "Radiant Faith Living with Ann Graham Lotz", "Living Life in God’s Illuminating Light with Ann Graham Lotz". Additionally, we quantify image quality of the concepts by computing FID score (denoted FIDConcept) between 10,000 images generated from the prompts before and after pruning for VM and TM samples.
@@
The results in Table˜10 show that the concepts associated with the memorized images suffer after mitigation with Wanda. We observe a significant drop from $\bm{A}_{\text{CLIP}}$ of around $0.37$ to $0.33$ for VM, which suggests that the generated images no longer follow the prompt. Moreover, the quality of the generated images degrades. FIDConcepts above 80 for VM and above 90 for TM samples indicates significant harm to the model, corroborated by the last row of Table˜8. Additionally, we provide qualitative results of damage to concepts in Section˜K.4.
@@
[…]
No Mitigation Verbatim $0.90\pm 0.04$ $0.37\pm 0.02$ N/A $1.00$
Template $0.17\pm 0.09$ $0.36\pm 0.02$ N/A $1.00$
Wanda + 10% pruned + [Uncaptioned image] Verbatim $0.40\pm 0.13$ $0.33\pm 0.02$ 80.70 $0.01$
Template $0.22\pm 0.08$ $0.33\pm 0.02$ 92.46 $0.01$
Appendix I Additional Details and Experiments on Adversarial Fine-Tuning
I.1 Algorithmic Description
Alg.˜2 provides an algorithmic overview of our adversarial fine-tuning mitigation method. As described in the main paper, we begin by generating images from memorized prompts using a mitigation technique that preserves alignment with the prompt while avoiding replication of training data. Alternatively, these images can be generated using a separate DM that has not been trained on the corresponding samples. To preserve the model’s general utility, a second set of non-memorized samples is incorporated during fine-tuning.
@@ I.1 Algorithmic Description
Algorithm 2 Fine-Tuning DM to Mitigate Memorization
  DM $\bm{\epsilon}_{\bm{\theta}}$
@@
   Surrogate images $\mathcal{D}_{\text{surrogate}}$ $\triangleright$ Images generated with active mitigation
@@
     ${\bm{y}}_{\text{adv}}\leftarrow\text{find\_adv\_embedding}({\bm{x}}_{\text{mem}},{\bm{p}}_{\text{mem}})$ $\triangleright$ Start from text embedding
@@
     ${\bm{y}}_{\text{adv}}\leftarrow\text{find\_adv\_embedding}({\bm{x}}_{\text{mem}},\text{random})$ $\triangleright$ Start from random embedding
@@
     $\bm{\epsilon}_{\text{surr}}\sim\mathcal{N}(0,\bm{I})$ $\triangleright$ Update with surrogate/adversarial sample
@@
     $\bm{\epsilon}_{\text{non-mem}}\sim\mathcal{N}(0,\bm{I})$ $\triangleright$ Update with non-memorized sample
@@
     $\mathcal{L}_{\text{total}}\leftarrow\mathcal{L}_{\text{adv}}+\mathcal{L}_{\text{non-mem}}$ $\triangleright$ Aggregate both losses
@@
I.2 Qualitative Examples of Surrogate Images
Figure 7: Examples of surrogate images used for adversarial fine-tuning. Multiple surrogate samples were generated for the memorized image (top left), preserving its semantic content without replicating the original. The images were created using NeMo to prevent exact replication.
I.3 Sensitivity Analysis
@@
Table˜11 presents results for using 25, 50, and 100 adversarial embedding optimization steps during training to craft the adversarial embeddings used in the mitigation loss $\mathcal{L}_{\mathit{Adv}}$. The number of steps used during training is indicated in the Setting column. We further evaluate the mitigation effect using the unchanged embeddings of the memorized training prompts (denoted by 0 in the Adv. Steps column), as well as the same embeddings optimized with 50 steps. Results are reported only for VM prompts, which the model has been fine-tuned on.
Table˜12 repeats the previous analysis, but explores the impact of starting the adversarial optimization from random embeddings instead of training prompt embeddings. We additionally compare the mitigation effect of adversarial embeddings crafted with 50 and 100 optimization steps, respectively.
@@
No Mitigation 0 $0.88\pm 0.06$ $1.0\pm 0.0$ $0.99\pm 0.01$ $0.32\pm 0.02$ $0.98$ $14.44$ $0.0060$
Training with 25 adv. steps 0 1 $0.16\pm 0.05$ $0.17\pm 0.06$ $5\pm 0.09 0.33±0$$3\pm 0.02$ $0.00$ $15.26$ $0.0058$
5 $0.15\pm 0.06$ $0.16\pm 0.07$ $0.40\pm 0.09$ $33\pm 0.02 0.0$00.00 $14.33$ $0.0052$
10 $0.14\pm 0.07$ $0.15\pm 0.07$ $0.43\pm 0.09$ $33\pm 0.02 0.0$00.00 $14.86$ $0.0052$
20 $0.14\pm 0.06$ $0.16\pm 0.06$ $0.49\pm 0.10$ $32\pm 0.02 0.0$00.00 $15.03$ $0.0047$
30 $0.12\pm 0.06$ $0.14\pm 0.05$ $0.52\pm 0.10$ $33\pm 0.01 0.0$00.00 $15.46$ $0.0049$
40 $0.13\pm 0.05$ $0.14\pm 0.05$ $0.54\pm 0.10$ $32\pm 0.02 0.0$00.00 $15.56$ $0.0047$
50 $0.12\pm 0.07$ $0.14\pm 0.07$ $0.54\pm 0.11$ $32\pm 0.02 0.0$00.00 $15.52$ $0.0047$
Training with 25 adv. steps 50 1 $0.57\pm 0.24$ $0.59\pm 0.23$ $0.64\pm 0.26$ $0.32\pm 0.01$ $0.37$ $15.26$ $0.0058$
5 $0.38\pm 0.18$ $0.39\pm 0.18$ $0.51\pm 0.12$ $0.32\pm 0.01$ $0.12$ $14.33$ $0.0052$
10 $0.38\pm 0.16$ $0.39\pm 0.19$ $0.56\pm 0.15$ $0.32\pm 0.02$ $0.11$ $14.86$ $0.0052$
20 $0.39\pm 0.18$ $0.40\pm 0.18$ $0.56\pm 0.13$ $0.32\pm 0.02$ $0.24$ $15.03$ $0.0047$
30 $0.32\pm 0.13$ $0.36\pm 0.16$ $0.51\pm 0.12$ $0.32\pm 0.02$ $0.15$ $15.46$ $0.0049$
40 $0.27\pm 0.11$ $0.29\pm 0.13$ $0.53\pm 0.10$ $32\pm 0.02 0.0$70.07 $15.56$ $0.0047$
50 $0.32\pm 0.15$ $0.34\pm 0.14$ $0.52\pm 0.09$ $0.33\pm 0.02$ $0.12$ $15.52$ $0.0047$
Training with 50 adv. steps 0 1 $0.14\pm 0.05$ $0.15\pm 0.05$ $3\pm 0.08 0.33±0$$3\pm 0.02$ $0.00$ $15.66$ $0.0062$
5 $0.15\pm 0.07$ $0.15\pm 0.07$ $5\pm 0.08 0.33±0$$3\pm 0.01$ $0.00$ $13.61$ $0.0047$
10 $0.13\pm 0.05$ $0.14\pm 0.06$ $7\pm 0.09 0.33±0$$3\pm 0.02$ $0.00$ $15.16$ $0.0049$
20 $0.12\pm 0.05$ $0.13\pm 0.05$ $0.44\pm 0.08$ $32\pm 0.02 0.0$00.00 $15.56$ $0.0051$
30 $0.14\pm 0.05$ $0.14\pm 0.05$ $0.45\pm 0.09$ $32\pm 0.02 0.0$00.00 $15.47$ $0.0053$
40 $0.12\pm 0.05$ $0.14\pm 0.06$ $0.50\pm 0.08$ $32\pm 0.01 0.0$00.00 $16.65$ $0.0055$
50 $0.12\pm 0.05$ $0.14\pm 0.06$ $0.52\pm 0.09$ $32\pm 0.01 0.0$00.00 $16.02$ $0.0055$
Training with 50 adv. steps 50 1 $0.64\pm 0.16$ $0.69\pm 0.16$ $0.75\pm 0.20$ $0.32\pm 0.01$ $0.17$ $15.66$ $0.0062$
5 $0.36\pm 0.14$ $0.38\pm 0.14$ $0.54\pm 0.10$ $30\pm 0.02 0.0$20.02 $13.61$ $0.0047$
10 $0.26\pm 0.15$ $0.27\pm 0.16$ $0.46\pm 0.10$ $30\pm 0.02 0.0$10.01 $15.16$ $0.0049$
20 $0.29\pm 0.13$ $0.30\pm 0.13$ $0.46\pm 0.07$ $30\pm 0.02 0.0$00.00 $15.56$ $0.0051$
30 $0.23\pm 0.11$ $0.28\pm 0.12$ $0.46\pm 0.10$ $30\pm 0.02 0.0$00.00 $15.47$ $0.0053$
40 $0.22\pm 0.12$ $0.25\pm 0.13$ $0.48\pm 0.09$ $31\pm 0.02 0.0$00.00 $16.65$ $0.0055$
50 $0.19\pm 0.10$ $0.21\pm 0.11$ $0.46\pm 0.06$ $31\pm 0.02 0.0$00.00 $16.02$ $0.0055$
Training with 100 adv. steps 0 1 $0.13\pm 0.04$ $0.14\pm 0.05$ $0.29\pm 0.07$ $32\pm 0.02 0.0$00.00 $15.32$ $0.0051$
5 $0.13\pm 0.05$ $0.14\pm 0.05$ $0.30\pm 0.06$ $32\pm 0.02 0.0$00.00 $14.36$ $0.0049$
10 $0.13\pm 0.05$ $0.14\pm 0.06$ $0.34\pm 0.09$ $32\pm 0.02 0.0$00.00 $15.56$ $0.0051$
20 $0.13\pm 0.05$ $0.13\pm 0.05$ $0.38\pm 0.08$ $32\pm 0.02 0.0$00.00 $15.47$ $0.0053$
30 $0.12\pm 0.05$ $0.14\pm 0.05$ $0.42\pm 0.07$ $32\pm 0.02 0.0$00.00 $15.34$ $0.0053$
40 $0.12\pm 0.04$ $0.13\pm 0.04$ $0.43\pm 0.09$ $32\pm 0.02 0.0$00.00 $16.23$ $0.0052$
50 $0.12\pm 0.04$ $0.13\pm 0.04$ $0.46\pm 0.09$ $32\pm 0.02 0.0$00.00 $17.52$ $0.0056$
Training with 100 adv. steps 50 1 $0.58\pm 0.13$ $0.61\pm 0.11$ $0.73\pm 0.13$ $31\pm 0.01 0.0$50.05 $15.32$ $0.0051$
5 $0.31\pm 0.10$ $0.32\pm 0.12$ $0.44\pm 0.07$ $30\pm 0.02 0.0$10.01 $14.36$ $0.0049$
10 $0.29\pm 0.11$ $0.30\pm 0.12$ $0.42\pm 0.08$ $29\pm 0.02 0.0$00.00 $15.56$ $0.0051$
20 $0.25\pm 0.11$ $0.27\pm 0.11$ $0.40\pm 0.07$ $29\pm 0.02 0.0$00.00 $15.47$ $0.0053$
30 $0.19\pm 0.09$ $0.20\pm 0.11$ $0.41\pm 0.08$ $30\pm 0.02 0.0$00.00 $15.34$ $0.0053$
40 $0.22\pm 0.10$ $0.24\pm 0.10$ $0.41\pm 0.06$ $30\pm 0.02 0.0$00.00 $16.23$ $0.0052$
50 $0.19\pm 0.10$ $0.20\pm 0.09$ $0.42\pm 0.07$ $30\pm 0.02 0.0$00.00 $17.52$ $0.0056$
[…]
No Mitigation 0 $0.88\pm 0.06$ $1.0\pm 0.0$ $0.99\pm 0.01$ $0.32\pm 0.02$ $1.00$ $14.44$ $0.0060$
Training with 25 adv. steps 50 1 $0.55\pm 0.23$ $0.58\pm 0.22$ $0.62\pm 0.24$ $0.32\pm 0.02$ $0.51$ $15.26$ $0.0058$
5 $0.49\pm 0.20$ $0.49\pm 0.22$ $0.53\pm 0.16$ $0.31\pm 0.02$ $0.27$ $14.33$ $0.0052$
10 $0.42\pm 0.18$ $0.45\pm 0.19$ $0.54\pm 0.14$ $0.31\pm 0.02$ $0.21$ $14.86$ $0.0052$
20 $0.45\pm 0.18$ $0.45\pm 0.20$ $0.57\pm 0.16$ $0.31\pm 0.02$ $0.24$ $15.03$ $0.0047$
30 $0.35\pm 0.16$ $0.39\pm 0.17$ $0.51\pm 0.15$ $0.31\pm 0.02$ $0.17$ $15.46$ $0.0049$
40 $0.28\pm 0.14$ $0.29\pm 0.15$ $0.52\pm 0.12$ $0.31\pm 0.02$ $0.10$ $15.56$ $0.0047$
50 $0.37\pm 0.17$ $0.39\pm 0.17$ $0.52\pm 0.11$ $0.32\pm 0.02$ $0.19$ $15.52$ $0.0047$
Training with 25 adv. steps 100 1 $9\pm 0.05 0.88±0$$8\pm 0.07$ $0.99\pm 0.01$ $0.32\pm 0.02$ $0.88$ $15.26$ $0.0058$
5 $0.82\pm 0.10$ $0.81\pm 0.12$ $0.89\pm 0.11$ $0.32\pm 0.02$ $0.67$ $14.33$ $0.0052$
10 $4\pm 0.15 0.77±0$$7\pm 0.16$ $0.85\pm 0.14$ $0.32\pm 0.01$ $0.55$ $14.86$ $0.0052$
20 $0.76\pm 0.14$ $0.79\pm 0.14$ $0.86\pm 0.13$ $0.32\pm 0.02$ $0.55$ $15.03$ $0.0047$
30 $6\pm 0.12 0.77±0$$7\pm 0.12$ $0.92\pm 0.08$ $0.32\pm 0.02$ $0.60$ $15.46$ $0.0049$
40 $0.68\pm 0.20$ $0.70\pm 0.23$ $0.63\pm 0.24$ $0.32\pm 0.02$ $0.49$ $15.56$ $0.0047$
50 $0.76\pm 0.12$ $0.80\pm 0.13$ $0.83\pm 0.16$ $0.32\pm 0.01$ $0.60$ $15.52$ $0.0047$
Training with 50 adv. steps 50 1 $0.64\pm 0.17$ $0.68\pm 0.16$ $0.72\pm 0.20$ $0.32\pm 0.01$ $0.44$ $15.66$ $0.0062$
5 $0.37\pm 0.12$ $0.39\pm 0.14$ $0.54\pm 0.11$ $30\pm 0.02 0.0$40.04 $13.61$ $0.0047$
10 $0.26\pm 0.15$ $0.27\pm 0.16$ $0.46\pm 0.11$ $30\pm 0.02 0.0$10.01 $15.16$ $0.0049$
20 $0.29\pm 0.12$ $0.30\pm 0.13$ $0.46\pm 0.07$ $30\pm 0.02 0.0$20.02 $15.56$ $0.0051$
30 $0.23\pm 0.11$ $0.28\pm 0.13$ $0.46\pm 0.10$ $30\pm 0.02 0.0$20.02 $15.47$ $0.0053$
40 $0.22\pm 0.12$ $0.25\pm 0.12$ $0.48\pm 0.09$ $31\pm 0.02 0.0$20.02 $16.65$ $0.0055$
50 $0.19\pm 0.10$ $0.21\pm 0.11$ $0.46\pm 0.06$ $31\pm 0.02 0.0$10.01 $16.02$ $0.0055$
Training with 50 adv. steps 100 1 $0.83\pm 0.08$ $0.85\pm 0.07$ $0.94\pm 0.06$ $0.32\pm 0.01$ $0.73$ $15.66$ $0.0062$
5 $0.54\pm 0.18$ $0.56\pm 0.17$ $0.67\pm 0.20$ $0.31\pm 0.01$ $0.29$ $13.61$ $0.0047$
10 $0.47\pm 0.20$ $0.45\pm 0.20$ $0.52\pm 0.17$ $0.31\pm 0.02$ $0.21$ $15.16$ $0.0049$
20 $0.46\pm 0.20$ $0.47\pm 0.20$ $0.61\pm 0.18$ $0.31\pm 0.02$ $0.20$ $15.56$ $0.0051$
30 $0.37\pm 0.19$ $0.38\pm 0.19$ $0.53\pm 0.15$ $0.31\pm 0.02$ $0.16$ $15.47$ $0.0053$
40 $0.40\pm 0.19$ $0.40\pm 0.21$ $0.55\pm 0.13$ $0.32\pm 0.02$ $0.18$ $16.65$ $0.0055$
50 $0.34\pm 0.17$ $0.34\pm 0.18$ $0.52\pm 0.10$ $0.32\pm 0.01$ $0.13$ $16.02$ $0.0055$
Training with 100 adv. steps 50 1 $0.59\pm 0.13$ $0.61\pm 0.10$ $0.72\pm 0.13$ $0.32\pm 0.01$ $0.31$ $15.32$ $0.0051$
5 $2\pm 0.11 0.33±0$$3\pm 0.12$ $0.44\pm 0.07$ $30\pm 0.02 0.0$20.02 $14.36$ $0.0049$
10 $0.30\pm 0.11$ $0.30\pm 0.12$ $0.42\pm 0.08$ $29\pm 0.02 0.0$10.01 $15.56$ $0.0051$
20 $0.25\pm 0.11$ $0.27\pm 0.12$ $0.40\pm 0.07$ $29\pm 0.02 0.0$00.00 $15.47$ $0.0053$
30 $0.19\pm 0.09$ $0.20\pm 0.11$ $0.41\pm 0.08$ $30\pm 0.02 0.0$20.02 $15.34$ $0.0053$
40 $0.22\pm 0.10$ $0.24\pm 0.11$ $0.41\pm 0.06$ $30\pm 0.02 0.0$00.00 $16.23$ $0.0052$
50 $0.19\pm 0.10$ $0.20\pm 0.09$ $0.42\pm 0.07$ $30\pm 0.02 0.0$00.00 $17.52$ $0.0056$
Training with 100 adv. steps 100 1 $0.77\pm 0.11$ $0.79\pm 0.10$ $0.86\pm 0.13$ $0.32\pm 0.01$ $0.64$ $15.32$ $0.0051$
5 $0.38\pm 0.11$ $0.37\pm 0.13$ $0.48\pm 0.10$ $30\pm 0.02 0.0$50.05 $14.36$ $0.0049$
10 $0.34\pm 0.13$ $0.36\pm 0.13$ $0.48\pm 0.07$ $30\pm 0.02 0.0$40.04 $15.56$ $0.0051$
20 $0.26\pm 0.13$ $0.29\pm 0.14$ $0.48\pm 0.10$ $29\pm 0.02 0.0$40.04 $15.47$ $0.0053$
30 $0.22\pm 0.11$ $0.23\pm 0.12$ $0.43\pm 0.08$ $30\pm 0.02 0.0$40.04 $15.34$ $0.0053$
40 $0.22\pm 0.12$ $0.24\pm 0.13$ $0.48\pm 0.09$ $30\pm 0.02 0.0$40.04 $16.23$ $0.0052$
50 $0.19\pm 0.10$ $0.22\pm 0.12$ $0.48\pm 0.09$ $31\pm 0.02 0.0$40.04 $17.52$ $0.0056$
I.4 Ablation Study
We perform an ablation study to evaluate the impact of each individual component of the adversarial fine-tuning procedure. Specifically, we compare the default setting, where adversarial embeddings optimized for 50 steps are used over three consecutive fine-tuning steps, with alternative configurations. In one variant, only a single update is performed per embedding (One Step). In another, the model is fine-tuned exclusively on samples from the surrogate set (No Utility Loss). We also assess a setting where only non-memorized samples are used during fine-tuning (No Mitigation Loss). All configurations are evaluated across different training epochs, both when using the original training prompts without any adversarial optimization, and when using adversarial embeddings optimized for 50 steps.
The results in Table˜13 show that fine-tuning the model for only a single step per adversarial embedding per epoch already achieves effective mitigation. This suggests that the fine-tuning process can be accelerated by reducing the number of updates per embedding. When the model is trained exclusively on surrogate samples—aiming to mitigate memorization without including any non-memorized samples to preserve utility—we observe strong mitigation, but at the cost of significantly degraded image quality, as reflected in the high FID and KID scores. Conversely, fine-tuning only on non-memorized samples while excluding surrogate samples helps maintain image quality but fails to provide sufficient mitigation against data replication.
[…]